OpenStudy (anonymous):

Which equation represents a quadratic function? (1) y = x + 2 (3) y = x 2 (2) y = ∙x + 2∙ (4) y = 2

OpenStudy (anonymous):

none of then - a quadratic equation will contain a squared variable like x^2

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

* quadratic function

OpenStudy (anonymous):

Well that's the only choices I have, It doesnt say none of these above.

OpenStudy (anonymous):

hhm - think so y = x^2 y = x^2 + 3x -2 are quadratics

OpenStudy (anonymous):

Which one from the multiple choice is it?

OpenStudy (anonymous):

- i'm going by the uk definition - maybe its different in other countries all the above options are linear aren't they?

OpenStudy (anonymous):

Yes but thats not of the options man

OpenStudy (anonymous):

I gave you the multiple choice above.

OpenStudy (anonymous):

ioh - is the no 3 y = x2 ( or x^2) ? if so thats the one

OpenStudy (anonymous):

- i thought maybe that was a typo!

OpenStudy (anonymous):

Thanks Dude

OpenStudy (anonymous):

yw man